OpenStudy (anonymous):

What is the slope that passes through points (8,1) and (-3,-6) on a graph

OpenStudy (anonymous):

Slope is usually represented by the letter m. Slope m is equal to (change in y) divided by (change in x). (-6-1) / (-3-8) = -7/(-11) = 7/11

OpenStudy (anonymous):

So the slope of the line that goes thru those 2 points is 7/11.